Analytic Geometry
2. For a line segment AB, one endpoint is A(-6, 2) and the other endpoint is B(1, -5). Find the coordinates of the midpoint M.
Thank you very much!!!!

Answer by stanbon(75887)

For a line segment AB, one endpoint is A(-6, 2) and the other endpoint is B(1, -5). Find the coordinates of the midpoint M.
--------------------
Midpoint coordinates are the average of the corresponding endpoint coordinates.
-----------------------
Let the midpoint be (x,y)
---
x = (-6+1)/2 = -5/2
y = (2 + 5)/2 = 7/2
==========================
midpoint = (-5/2,7/2)
=======================
Cheers,
Stan H.